OpenStudy (anonymous):

can someone please walk me through how to solve this problem? I followed the steps in the book but got every single one of them wrong. Root(-17) * root(-51)

OpenStudy (anonymous):

Root(-17) * root(-51)=root((-17)*(-51))=root(867)

OpenStudy (anonymous):

what are the step involved in getting the answer? That's the problem I'm having

OpenStudy (anonymous):

huh i am stumped

OpenStudy (kira_yamato):

We know that √–1 = i So i^2 = -1 √-17 * √-51 = i√17 * i√51 = i^2 * √(17*51) = - √867

OpenStudy (thomas5267):

\[ \begin{align*} &\sqrt{-17}\times\sqrt{-51} \\ =&\sqrt{-17\times-51} \\ =&\sqrt{867} \end{align*} \]